the diameter of a can is 6cm and the height of the can is 8cm. How many cm cubed will the can hold? tell how you get the answer

Answer by checkley77(12844)

V=BH
THE BASE IS A CIRCLE.
AREA OF A CIRCLE=PI*R^2
V=PI*R^2*H
V=3.14(6/2)^2*8
V=3.14*3^2*8
V=3.14*9*8
V=226.08 CM^3 ANS.